OpenStudy (anonymous):

simplify n^6 · n^5 · n^4 ÷ n^3 · n^2 ÷ n help me please!!

OpenStudy (anonymous):

n(6 · 5 · 4 ÷ 3 · 2 ÷ 1) n(30 · 4 ÷ 3 · 2 ÷ 1) n(120 ÷ 3 · 2 ÷ 1) n(40 · 2 ÷ 1) n(80 ÷ 1) n(80) 80n <------------------------------

OpenStudy (anonymous):

i think thats wrong nikka

OpenStudy (lalaly):

no thats not right \[\huge{\frac{n^6n^5n^4}{\frac{n^3n^2}{n}}}=\frac{n^6n^5n^4}{n^4}=\frac{n^{15}}{n^4}=n^{11}\]

OpenStudy (anonymous):

OOPS yep I see what I did wrong